I need step by step calculation!

Ohms Law A resistor rated at 250 k Ω is connected across two D cell batteries (each 1 .50 V) in series, with a total voltage of 3.00 V. The manufacturer advertises that their resistors are within 5% of the rated value. What are the possible minimum current and maximum current through the resistor?
Answer #1

R = 250 kΩ

V = 3.0 V

Rmin = 250*0.95 = 237.5 kΩ

Rmax = 250*1.05 = 262.5 kΩ

So Imax = V/Rmin = 3/237.5 = 12.6 uA

and Imin = V/Rmax = 3/262.5 = 11.4 uA
Answer #2

SOLUTION :


R = 250 kΩ 


Variation possible I resistance value = 5% I.e. - 5% to + 5% .


Hence,


R minimum = 250 * (1 - 0.05) = 237.5 kΩ 


R maximum = 250* (1 + 0.05) = 262.5 kΩ


Voltage applied = 2 * 1,50 = 3.00 V


So, 


Minimum current = 3.00 / (262.5 *10^3) = 11.43 * 10^(-6) Amp = 11.43 µA (ANSWER).


Maximum current = 3.00 / (237.5 *10^3) = 12.633 * 10^(-6) Amp = 12.63 µA (ANSWER).

answered by: Tulsiram Garg